# Deep Height Decoupling for Precise Vision-based 3D Occupancy Prediction

### v1

@@ Abstract
The task of vision-based 3D occupancy prediction aims to reconstruct 3D geometry and estimate its semantic classes from 2D color images, where the 2D-to-3D view transformation is an indispensable step. Most previous methods conduct forward projection, such as BEVPooling and VoxelPooling, both of which map the 2D image features into 3D grids. However, the current grid representing features within a certain height range usually introduces many confusing features that belong to other height ranges. To address this challenge, we present Deep Height Decoupling (DHD), a novel framework that incorporates explicit height prior to filter out the confusing features. Specifically, DHD first predicts height maps via explicit supervision. Based on the height distribution statistics, DHD designs Mask Guided Height Sampling (MGHS) to adaptively decoupled the height map into multiple binary masks. MGHS projects the 2D image features into multiple subspaces, where each grid contains features within reasonable height ranges. Finally, a Synergistic Feature Aggregation (SFA) module is deployed to enhance the feature representation through channel and spatial affinities, enabling further occupancy refinement. On the popular Occ3D-nuScenes benchmark, our method achieves state-of-the-art performance even with minimal input frames. Code is available at https://github.com/yanzq95/DHD.
@@ I Introduction
Understanding the 3D geometry and semantic information of surrounding environment is crucial for autonomous driving. In recent years, the rapid development of camera-based algorithms [1, 2, 3, 4, 5, 6] has significantly advanced outdoor 3D scene understanding, leading to increased attention on the vision-based 3D occupancy prediction task [7, 8, 9, 10, 11, 12].
@@
To address these challenges, we propose a novel framework called deep height decoupling (DHD). For the first time, DHD introduces explicit height prior to decouple the 3D features and filter out the redundant parts. Specifically, DHD first generates height maps through LiDAR supervisory signals, capturing the height distribution of the scene. Then, to achieve precise and effective 2D-to-3D view transformation, DHD designs a mask guided height sampling (MGHS) module to sample features across varying heights. Based on reasonable height distribution statistics, MGHS decouples the height maps into multiple height masks, each corresponding to a distinct height range. These masks are applied to filter out the 2D features for height-aware feature sampling. Finally, the masked 2D features are projected into multiple 3D subspaces, ensuring that the 3D features within each grid fall within reasonable height ranges. Additionally, DHD introduces a synergistic feature aggregation (SFA) module, which enhances the dual 3D feature representations by leveraging both channel and spatial affinities, contributing to further occupancy refinement.
@@
  • DHD achieves superior performance with minimal input cost. Codes are released for peer research.
II Relate Work
@@ II-A Vision-Based 3D Occupancy Prediction
Recently, vision-based 3D occupancy prediction [9, 8, 7, 10] has garnered increasing attention, including both supervised and unsupervised methods. For supervised methods [1, 19, 2, 7], MonoScene [19] is the pioneering approach for monocular occupancy prediction. BEVDet [1] employs Lift-Splat-Shoot (LSS) [20] to project 2D image features into 3D space. BEVDet4D [21] further explores the temporal fusion strategy by fusing features from the previous frames. Besides, several transformer-based methods [2, 7, 22, 23] have been proposed. For example, BEVformer [2] utilizes spatiotemporal transformers to construct BEV features. VoxFormer [7] introduces a sparse-to-dense transformer framework for 3D semantic scene completion. For unsupervised methods [24, 25, 26, 5], SelfOcc [24] and OccNeRF [25] are two representative works that employ volume rendering and photometric consistency to generate self-supervised signals. Different from previous methods that directly extract 3D features, we introduce height prior to decouple height, effectively capturing finer details across different height ranges.
@@
To transform 2D image features into 3D space, numerous view transform methods have been proposed. LSS [20] explicitly predicts the depth distribution and elevates 2D image features into 3D space. Most recently, a few methods [13, 14, 27] emphasize the significance of accurate depth estimation in view transformation. For example, BEVDepth [13] encodes camera intrinsics and extrinsics into a depth refinement module for 3D object detection. BEVStereo [14] introduces an effective temporal stereo technique to enhance depth estimation. Additionally, some researches [28, 29, 30, 31] focus on optimizing the projection stage. For instance, SparseFusion [29] argues that projecting all virtual points into the BEV space is unnecessary. Similarly, SA-BEV [28] utilizes SA-BEVPool to project only foreground image features. Differently, we employ height masks to selectively project features, enhancing the accuracy of spatial representation.
@@ III-A Overview
Figure 4: (a) We divide the height into three spaces to differentiate features across heights and list the proportion of each class below. (b) The distribution of various categories across different heights, with the bar chart presenting statistical values within each subspace.
@@ III-B HeightNet
Furthermore, we propose using ground-truth $\boldsymbol{H}_{\mathrm{gt}}$ from LiDAR points to supervise the learning of $\boldsymbol{H}_{\mathrm{pred}}$, facilitating the accurate height estimation. To obtain $\boldsymbol{H}_{\mathrm{gt}}$, given a 3D point $\boldsymbol{p}_{\mathrm{l}}=[x_{\mathrm{l}},y_{\mathrm{l}},z_{\mathrm{l}},1]^{\top}$ in the LiDAR coordinate system, we calculate its translate position $\boldsymbol{p}_{\mathrm{e}}=[x_{\mathrm{e}},y_{\mathrm{e}},z_{\mathrm{e}},1]^{\top}$ in the ego coordinate system:
[…]
where $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$ and $\boldsymbol{t}\in\mathbb{R}^{3}$ denote the rotation and translation matrices, respectively, for transforming from LiDAR coordinate to the ego coordinate.
@@ III-C Mask Guided Height Sampling
$\blacksquare$ cons. veh. $\uparrow$
@@ III-D Synergistic Feature Aggregation
The key idea of the aggregation module is to select and construct the most relevant features for the occupancy prediction. Our SFA is shown in Fig. 7. The depth-based feature $\boldsymbol{F}_{\mathrm{db}}$ and height-refined feature $\boldsymbol{F}_{\mathrm{hr}}$ are first input into the channel stage, producing the affinity vector $\boldsymbol{a}_{1}$:
@@ III-E Training Loss
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are hyper-parameters. Noted that in DHD-S, depth supervision is not utilized.
@@ IV-C Comparison with the State-of-the-Art
Visual Comparison. Fig. 8 presents the visual comparison of DHD and FlashOcc [16] configured with a resolution of $512\times 1408$. Notably, our method can predict more accurate occupancy results. For example, compared to FlashOcc [16], the first row demonstrates that our DHD performs better on drive. surf. and terrain, both of which are primarily concentrated at lower heights. This improvement highlights DHD’s ability to effectively handle lower-level features. The second and third rows show that our model excels in differentiating objects within similar height ranges, such as truck and trailer. Additionally, DHD accurately estimate the instance of pedestrian and vegetation. These visual results demonstrate that our approach achieves a more accurate understanding of the 3D scene.
@@ V CONCLUSIONS
In this paper, we proposed the novel deep height decoupling (DHD) framework for vision-based 3D occupancy prediction. For the first time, DHD introduced the explicit height prior to filter the confusing occupancy features. Specifically, DHD estimated the height map and decoupled it into multiple masks based on very solid height distribution statistics, including heatmap visualization, cumulative distribution function curve, and entropy calculation. Then the mask guided height sampling was designed to realize the more accurate 2D-to-3D view transformation. In addition, at the end of the model, the two-stage synergistic feature aggregation is introduced to enhance the feature representation using channel and spatial affinities. Owing to these designs, DHD achieved state-of-the-art performance even with minimal input frames on Occ3D-nuScenes benchmark.

### v3

@@ Abstract
The task of vision-based 3D occupancy prediction aims to reconstruct 3D geometry and estimate its semantic classes from 2D color images, where the 2D-to-3D view transformation is an indispensable step. Most previous methods conduct forward projection, such as BEVPooling and VoxelPooling, both of which map the 2D image features into 3D grids. However, the current grid representing features within a certain height range usually introduces many confusing features that belong to other height ranges. To address this challenge, we present Deep Height Decoupling (DHD), a novel framework that incorporates explicit height prior to filter out the confusing features. Specifically, DHD first predicts height maps via explicit supervision. Based on the height distribution statistics, DHD designs Mask Guided Height Sampling (MGHS) to adaptively decouple the height map into multiple binary masks. MGHS projects the 2D image features into multiple subspaces, where each grid contains features within reasonable height ranges. Finally, a Synergistic Feature Aggregation (SFA) module is deployed to enhance the feature representation through channel and spatial affinities, enabling further occupancy refinement. On the popular Occ3D-nuScenes benchmark, our method achieves state-of-the-art performance even with minimal input frames. Source code is released at https://github.com/yanzq95/DHD.
@@ I Introduction
Understanding the 3D geometry and semantic information of surrounding environment is crucial for autonomous driving. In recent years, the rapid development of camera-based algorithms has significantly advanced outdoor 3D scene understanding [1, 2, 3, 4, 5, 6, 7, 8, 9], leading to increased attention on the vision-based 3D occupancy prediction task [10, 11, 12, 13, 14, 15].
@@
To address these challenges, we propose a novel framework called deep height decoupling (DHD). For the first time, DHD introduces explicit height prior to decouple the 3D features and filter out the redundant parts. Specifically, DHD first generates height maps through LiDAR supervisory signals, capturing the height distribution of the scene. Then, to achieve precise and effective 2D-to-3D view transformation, DHD designs a mask guided height sampling (MGHS) module to sample features across varying heights. Based on reasonable height distribution statistics, MGHS decouples the height maps into multiple height masks, each corresponding to a distinct height range. These masks are applied to filter out the 2D features for height-aware feature sampling. Finally, the masked 2D features are projected into multiple 3D subspaces, ensuring that the 3D features within each grid fall within reasonable height ranges. Additionally, DHD introduces a synergistic feature aggregation (SFA) module, which enhances the dual feature representations by leveraging both channel and spatial affinities, contributing to further occupancy refinement.
@@
  • DHD achieves superior performance with minimal input cost [^1] Using only one history frame in temporal fusion.. Codes are released for peer research.
II Related Work
@@ II-A Vision-Based 3D Occupancy Prediction
Recently, vision-based 3D occupancy prediction [12, 11, 13] has garnered increasing attention, including both supervised and unsupervised methods. For supervised methods [1, 22, 2, 10], MonoScene [22] is the pioneering approach for monocular occupancy prediction. BEVDet [1] employs Lift-Splat-Shoot (LSS) [23] to project 2D image features into 3D space. BEVDet4D [24] further explores the temporal fusion strategy by fusing features from the previous frames. Besides, several transformer-based methods [2, 10, 25] have been proposed. For example, BEVformer [2] utilizes spatiotemporal transformers to construct BEV features. VoxFormer [10] introduces a sparse-to-dense transformer framework for 3D semantic scene completion. For unsupervised methods [26, 27, 28], SelfOcc [26] and OccNeRF [27] are two representative works that employ volume rendering and photometric consistency to generate self-supervised signals. Different from previous methods that directly extract 3D features, we introduce height prior to decouple height, effectively capturing finer details across different height ranges.
@@
Many existing methods leverage depth information [29, 30, 31] to perform 2D-to-3D view transformation. LSS [23] explicitly predicts the depth distribution and elevates 2D image features into 3D space. Most recently, a few methods [16, 17, 32] emphasize the significance of accurate depth estimation in view transformation. For example, BEVDepth [16] encodes camera intrinsics and extrinsics into a depth refinement module for 3D object detection. BEVStereo [17] introduces an effective temporal stereo technique to enhance depth estimation. Additionally, some researches [33, 34, 35, 36] focus on optimizing the projection stage. For instance, SparseFusion [34] argues that projecting all virtual points into the BEV space is unnecessary. Similarly, SA-BEV [33] utilizes SA-BEVPool to project only foreground image features. Differently, we employ height masks to selectively project features, enhancing the accuracy of spatial representation.
@@ III-A Overview
Figure 4: (a) We decouple height into three intervals to differentiate features across heights and list the proportion of each class below. (b) The distribution of various classes across different heights, with the bar chart presenting statistical values within each interval.
@@ III-B HeightNet
Furthermore, we propose using ground-truth $\boldsymbol{H}_{\mathrm{gt}}$ from LiDAR points to supervise the learning of $\boldsymbol{H}_{\mathrm{pred}}$, facilitating the accurate height estimation. To obtain $\boldsymbol{H}_{\mathrm{gt}}$, given a 3D point $\boldsymbol{p}_{\mathrm{l}}=[x_{\mathrm{l}},y_{\mathrm{l}},z_{\mathrm{l}},1]^{\top}$ in the LiDAR coordinate system, we calculate its translate position $\boldsymbol{p}_{\mathrm{e}}=[x_{\mathrm{e}},y_{\mathrm{e}},z_{\mathrm{e}},1]^{\top}$ in the ego coordinate system.
@@ III-C Mask Guided Height Sampling
$\blacksquare$ const. veh. $\uparrow$
@@ III-D Synergistic Feature Aggregation
The key idea of the aggregation module is to select and construct the most relevant features for the occupancy prediction through a two-stage attention mechanism. Our SFA is shown in Fig. 7. The depth-based feature $\boldsymbol{F}_{\mathrm{db}}$ and height-refined feature $\boldsymbol{F}_{\mathrm{hr}}$ are first input into the channel stage, producing the affinity vector $\boldsymbol{a}_{1}$:
@@ III-E Training Loss
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are hyper-parameters. Note that in DHD-S, depth supervision is not utilized.
@@ IV-C Comparison with the State-of-the-Art
Visual Comparison. Fig. 8 presents the visual comparison of DHD and FlashOcc [19] configured with a resolution of $512\times 1408$. Notably, our method can predict more accurate occupancy results. For example, compared to FlashOcc [19], the first row demonstrates that our DHD performs better on drive. surf. and terrain, both of which are primarily concentrated at lower heights. This improvement highlights DHD’s ability to effectively handle lower-level features. The second and third rows show that our model excels in differentiating objects within similar height ranges, such as truck and trailer. Additionally, DHD accurately estimate the instance of pedestrian and vegetation. These visual results demonstrate that our approach achieves a more precise and comprehensive understanding of the 3D scene.
@@ V CONCLUSIONS
In this paper, we proposed the novel deep height decoupling (DHD) framework for vision-based 3D occupancy prediction. For the first time, DHD introduced the explicit height prior to filter the confusing occupancy features. Specifically, DHD estimated the height map and decoupled it into multiple binary masks based on reliable height distribution statistics, including heatmap visualization, cumulative distribution function curve, and entropy calculation. Then the mask guided height sampling was designed to realize the more accurate 2D-to-3D view transformation. In addition, at the end of the model, the two-stage synergistic feature aggregation is introduced to enhance the feature representation using channel and spatial affinities. Owing to these designs, DHD achieved state-of-the-art performance even with minimal input frames on Occ3D-nuScenes benchmark.
ACKNOWLEDGMENT
This work was supported by the National Science Fund of China under Grant Nos. U24A20330 and 62361166670.